OpenStudy (anonymous):

(x-3)(x-3)=27 solve using an appropriate technique??? help please

OpenStudy (anonymous):

x^2-3x-3x+9=27 after solving lhs x^2-6x+9=27 x^2-6x-18=0 solve it further by yourself

OpenStudy (anonymous):

\[(x-3)^{2}=27\] \[x-3=\pm \sqrt{27}\] \[x-3=\sqrt{27} \] and \[x-3=-\sqrt{27}\]\[x=\sqrt{27}+3\] and \[x=-\sqrt{27}+3\]

OpenStudy (anonymous):

as \[\sqrt{27}=3\sqrt{3}\] therfore,\[x=3\sqrt{3}+3 \] and \[x=-3\sqrt{3}+3\]